OpenStudy (study.geek.123):

How do I find the x-intercepts of y=(x+2)^2+1

OpenStudy (erak):

Are you using imaginary numbers? Cause if not, there are none

OpenStudy (study.geek.123):

yes @Erak i am using imaginary numbers

OpenStudy (erak):

oh god i'm not the best but would turning this into standard form sound familiar?

OpenStudy (erak):

\[y = x^2 + 4x +5\]quadratic formula\[\frac{ -4 \pm \sqrt{16-20} }{ 2 }\]\[\frac{ -4 }{ 2 } \pm \frac{ \sqrt{4}i }{ 2 }\]\[-2 \pm i\]

OpenStudy (erak):

That's what I've learned just now, so hopefully somebody swings by to confirm

OpenStudy (study.geek.123):

thank you